Fermi National Laboratory


Accelerator Update for August 5, 2005 - August 8, 2005

Friday August 5
The day shift began with Operations monitoring store 4318 and the Recycler stash, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using colliding beam, with the Antiproton Source (Pbar) stacking, and with MiniBooNE, NuMI and the Meson MCenter experiments taking beam.

A Pbar expert began a study period at 9:51 AM.

Operations turned off the Linac beam at 1:01 PM so a Preaccelerator expert could access the I- Source to investigate the motor generator brushes. Power supply experts began work on a NuMI power supply (HV101) and techs began work on a Linac RF station (LRF5). The Preaccelerator expert was out of access by 1:14 PM.

The Linac techs completed their LRF5 work at 1:23 PM.

The power supply expert finished his work on HV101 at1:46 PM.

Operations established beam to NuMI at 2:02 PM.

At 5:59 PM, a Booster vacuum burst caused vacuum valves to close and when the valves closed a Booster power supply (GMPS) tripped off, as well as the 400 MeV area chopper.

At 6:54 PM, Booster RF stations 1, 3, 4, and 6 all tripped off. They reset okay.

MTest experiments said at 11:11 PM that they were done with the beam.

Saturday August 6
The midnight shift began with Operations monitoring store 4318 and the Recycler stash, with Pbar stacking and conducting studies, and with MiniBooNE, NuMI, MTest, and MCenter taking beam.

At 12:37 AM, NuMI tripped off for the first of seven times in four hours. It reset each time.

Operations terminated store 4318 at 3:02 AM and then put the Tevatron (TeV) into shot setup at 3:15 AM.

Operations established combination store 4319 at 5:46 AM with an initial luminosity of 117.34E30. Pbar resumed stacking.

Operations established beam to MCenter at 6:20 AM.

An Electron Cooling expert began a study period at 6:37 AM and finished at 8:17 AM. Then a Recycler expert began a study period.

At 9:09 AM, Operations reported that the NuMI autotune program kept stopping. Operators rebooted various systems and managed to get the autotune back on by 9:43 AM.

The Recycler study ended at 10:02 AM.

Operations put the Recycler into shot setup at 10:24 AM to stash antiprotons.

The Recycler stash ended and Pbar resumed stacking at 11:43 AM.

At 12:01 PM, a NuMI VME crate tripped off and took autotune down with it. Operators couldn't reset the crate and so called in experts.

The Recycler began stashing antiprotons at 4:38 PM. The stash ended and Pbar resumed stacking at 5:09 PM.

At 8:30 PM, BRF19 tripped off and caused all the other stations to begin sparking. It took a while to reset. It tripped off again at 9:57 PM. An operator called in an expert who had BRF19 back on by 10:47 PM.

Sunday, August 7
The midnight shift began with Operations monitoring store 4319 and the stash, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

At 1:18 AM, Operations reported that they couldn't keep BRF19 on and that every time it tripped off all the other RF stations would spark and some would trip off. Operations consulted experts.

At 2:46 AM, BRF19 tripped off due to a ground fault indication and would not stay on. An expert investigated.

At 3:13 AM, Operations turned off MiniBooNE and lowered NuMI's beam intensity while booster experts investigated. A few minutes later, operators turned off NuMI also.

Operations reestablished beam to MiniBooNE at 4:42 AM.

Operations reestablished low intensity beam to NuMI at 5:56 AM, but excessive sparking caused operators to stop the beam just eight minutes later.

Operations turned on the NuMI beam at 7:16 AM.

Operations terminated store 4319 at 12:08 PM and put the TeV back into shot setup at 12:25 PM.

After the final protons were loaded into the TeV, Operations began preparing the Booster at 1:45 PM for an access.

Operations established combination store 4320 at 2:38 PM with an initial luminosity of 112.2E30.

RF techs began working on BRF13 at 2:51 PM.

At 4:05 PM, a 480-volt distribution breaker tripped off while trying to check out BRF13. The Duty Electrician and a High Voltage expert closed in the breaker. There were several problems with this breaker tripping during this time period. At 6:17 PM, the Duty Electrician and HV Expert discovered that an interlock trip unit had failed.

Operations reported at 11:01 PM that the RF techs working in Booster were finishing things up and that the 480-volt interlock trip unit had been replaced.

Operations reported at 11:48 PM that the Booster RF was up and running.

Monday August 8
The midnight shift began with Operations monitoring store 4320, and with Booster holding off all downstream beam.

At 12:25 AM, Operations reported more Booster problems and called in experts.

Operations established beam to MiniBooNE at 1:18 AM.

Operations established beam to NuMI and Pbar resumed stacking at 1:25 AM.

Operations established beam to MCenter at 1:45 AM.

At 1:57 AM, Operations reported that BRF16 was tripping off repeatedly and would not stay on for more than a few minutes. Operators bypassed the station.

At 2:53 AM, Operations reported several trips of NuMI due to a MI kicker problem. This was also causing a stacking problem. At 3:17 AM, Operations reported that NuMI was getting beam, but that stacking has stopped.

At 5:04 AM, Operations reported that KPS5S wasn't charging due to a NIM crate timing problem. An operator called in experts.

Operators followed an expert's suggestion and bypassed the NIM timing problem. Pbar resumed stacking at 5:22 AM.

Recycler began shot setup at 6:39 AM to stash antiprotons.

Operators turned on BRF16 at 6:40 AM.

Experts reported at 6:51 AM that they had replaced the bad MI NIM crate.

The Recycler stash ended and Pbar resumed stacking at 7:38 AM.

MCenter resumed taking beam at 7:49 AM.

The Plan for Monday
The plan for today is to keep store 4320 in until this evening. Feeder 47 will be energized after the store is terminated, but no load will be put on it. A new store will be established with what antiprotons are available. The Run Coordinator will allow Booster, MI, and Recycler accesses tomorrow primarily to work on the Booster RF. Other jobs will be tacked on during this Booster downtime.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 8/8/2005   email Fermilab